10. (20 points) A horizontal force of N used to push a 40-kg box on a rough. horizontal surface through a distance of 6 m. If the box 150 is a constant speed, find: moves at a) the work done by the 150 N force b) the work done by friction c the coefficient of kinetic friction
Answer #1

a)work done by 150N force=F.d=150*6=900 J

b)since it is moving with constant speed so acceleration=0

By newtons law

F-f=0

F=f

f=150 N

work done by friction force=f.d=150*6=900 J

c)F-f=0

150-\muN=0

\mu=150/N

\mu=150/(40*9.8)=.38